Grade 12th passAlgebra

If G is the G.M between two positive numbers a and b, show that1/G²-a²+1/G²-b²=1/G².

Sumitkumarsingh
8 Years agoGrade 12th pass

2 Answers

Arun
ApprovedApproved Tutor Answer8 Years ago
Dear student
 
G is the G.M. between a & b.
G² = ab
 
Then LHS
1/(G²- a²)  +  1/(G²-b²)
= 1/ab- a²  +  1/ab - b²
= 1/a(b-a)  -  1/b(b-a)
= (b -a)/ (ab) *(b-a)
= 1/ab
= 1/G²
= RHS
Hence proved.
 
Regards
Arun (askIITians forum expert)
Sneha Khanal
4 Years ago
If A be the arithmetic vmean and H, the H.M between two quantities a and b, show that a-A/a-H * b-A/b-H = A/H